Komiko Tanaka invests $13,500 in LymaBean, Inc. LymaBean does not pay any dividends. Komiko projects that her investment will generate a 10 percent before-tax rate of return. She plans to invest for the long term. a. How much cash will Komiko retain, after-taxes, if she holds the investment for 5 years and then she sells it when the long-term capital gains rate is 15 percent? (Do not round your intermediate calculations. Round your final answer to the nearest whole dollar amount.) Cash retained b. What is Komikos after-tax rate of return on her investment in part (a)? (Do not round your intermediate calculations. Round your percentage answer to 2 decimal places.) After-tax rate of return percent c. How much cash will Komiko retain, after-taxes, if she holds the investment for 5 years and then sells when the long-term capital gains rate is 25 percent? (Do not round your intermediate calculations. Round your final answer to the nearest whole dollar amount.) Cash retainedd. What is Komikos after-tax rate of return on her investment in part (c)? (Do not round your intermediate calculations. Round your percentage answer to 2 decimal places.) After-tax rate of return percent e. How much cash will Komiko retain, after taxes, if she holds the investment for 15 years and then she sells when the long-term capital gains rate is 15 percent? (Do not round your intermediate calculations. Round your final answer to the nearest whole dollar amount.) Cash retained f. What is Komikos after-tax rate of return on her investment in part (e)? (Do not round your intermediate calculations. Round your percentage answer to 2 decimal places.) After-tax rate of return percent
Answer #1

a) Amount after 5 years=13500*(1+10%)^5
=21741.89
Capital gain=21741.89-13500=8241.89
Capital gain tax=8241.99*15%=1236.28
Cash retained=21741.89-1236.28=20506

b)After tax return=[(final/initial)^(1/5)]-1
=[(20506/13500)^(1/5)]-1
=8.72%

c) Amount after 5 years=13500*(1+10%)^5
=21741.89
Capital gain=21741.89-13500=8241.89
Capital gain tax=8241.99*25%=2060.47
Cash retained=21741.89-2060.47=19681

d)After tax return=[(final/initial)^(1/5)]-1
=[(19681/13500)^(1/5)]-1
=7.83%

e) Amount after 5 years=13500*(1+10%)^15
=56392.85
Capital gain=56392.85-13500=42892.85
Capital gain tax=42892.85*15%=6433.93
Cash retained=56392.85-6433.93=49959

f)After tax return=[(final/initial)^(1/5)]-1
=[(49959/13500)^(1/15)]-1
=9.12%